# Out-of-bag prediction balls for random forests in metric spaces

### v1

@@ Abstract
Statistical methods for metric spaces provide a general and versatile framework for analyzing complex data types. We introduce a novel approach for constructing confidence regions around new predictions from any bagged regression algorithm with metric-space-valued responses. This includes the recent extensions of random forests for metric responses: Fréchet random forests (Capitaine et al., 2024), random forest weighted local constant Fréchet regression (Qiu et al., 2024), and metric random forests (Bulté and Sørensen, 2024). Our prediction regions leverage out‐of‐bag observations generated during a single forest training, employing the entire data set for both prediction and uncertainty quantification. We establish asymptotic guarantees of out-of-bag prediction balls for four coverage types under certain regularity conditions. Moreover, we demonstrate the superior stability and smaller radius of out-of-bag balls compared to split-conformal methods through extensive numerical experiments where the response lies on the Euclidean space, sphere, hyperboloid, and space of positive definite matrices. A real data application illustrates the potential of the confidence regions for quantifying the uncertainty in the study of solar dynamics and the use of data-driven non-isotropic distances on the sphere.
@@
In Section A, Theorems 4.1 and 4.2 and Corollaries 4.1 and 4.2 are proved, relying on four technical lemmas. Section B shows numerical experiments to validate the Fréchet means of Wishart-distributed random variables under the AI and LC distances, given in (15) and (16), respectively. Section C contains a comprehensive list of figures and tables from Section 5. In Section D, we reapproach the derivations by Jensen, (1981, pp. 196–197) to clarify the connection of the HvMF and vMF distributions, and to detail the derivation and computation of the normalizing constant.
@@ 1 Introduction
As the complexity of data structures increases, traditional statistical methods that rely on vector space properties face significant limitations. Specialized methods have been developed for specific data types, including functional data (Li et al.,, 2022), directional data (Mardia and Jupp,, 1999), compositional data (Aitchison,, 1986), shape data (Dryden and Mardia,, 2016), and covariance matrices (Dryden et al.,, 2009). Random variables from these complex data types are studied by (Marron and Dryden,, 2021) under the terminology of random objects. A framework that has gained momentum in recent years is to consider random objects as elements in general metric spaces. While exploiting the structure of specific spaces can enhance the flexibility and tractability of statistical modeling, the generality of metric space analysis offers an unparalleled versatile framework that bypasses the need for algebraic structure or local Euclideanity, relying solely on the existence of a distance function.
@@
Key statistical concepts must be adapted to metric spaces. Fréchet, (1948) generalized the mean through square distance minimization. Extensions of regression models for non-Euclidean data have primarily focused on responses defined in Riemannian manifolds (e.g., Fletcher,, 2013; Kim et al.,, 2014). A foundational nonparametric regression estimator for metric output is the Nadaraya–Watson estimator, which uses local-constant weights determined by a kernel function and a bandwidth parameter. The Nadaraya–Watson estimator was extended by Hein, (2009) to the case when both the response and the predictors are random objects using a (generalized) Fréchet mean. Building on the notion of Fréchet mean, Petersen and Müller, (2019) introduced Fréchet regression, generalizing the standard Euclidean regression to random objects by estimating the conditional Fréchet mean. Under this setting, in ibid, global Fréchet regression was proposed as a parametric model that generalizes linear least squares regression for predictors in $\mathbb{R}^{p}$ and a metric response. Also in ibid, local Fréchet regression was introduced as a nonparametric approach that generalizes local linear estimation for random object responses with Euclidean predictors.
@@
Random Forests (RFs; Breiman,, 2001) is a popular nonparametric regression method widely recognized for its flexibility and predictive performance, being able to model complex, nonlinear relationships that other methods may struggle with, especially with high-dimensional data (Bühlmann and Yu,, 2002; Hastie et al.,, 2009; Varian,, 2014). In recent years, several authors have worked on adaptations of RFs to metric spaces under the framework of Fréchet regression. Capitaine et al., (2024) (first appeared as Capitaine et al.,, 2019) developed Fréchet Random Forests (FRFs), which adapt RFs for metric data by incorporating the Fréchet mean in the splitting criterion and the aggregation of trees. Alternatively, following the ideas of Meinshausen, (2006), Qiu et al., (2024) leveraged the weights generated in the building of each tree to estimate the response as a weighted average of the training responses. Bulté and Sørensen, (2024) introduced Fréchet medoids into the splitting criterion of the trees. Fréchet medoids serve as consistent estimators of the Fréchet mean and offer a computationally simpler alternative in settings where no closed-form expression is available and optimization over a manifold is typically employed.
@@
Despite their predictive power, RFs have historically lacked robust statistical inference tools, particularly in the estimation of the uncertainty in a RF prediction. For Euclidean data, Meinshausen, (2006) introduced quantile regression forests, which estimate conditional quantiles and naturally lead to prediction intervals. Zhang et al., (2020) proposed RF prediction intervals using out-of-bag observations, proving their asymptotic validity. This approach is conceptually related to conformal inference (Vovk et al.,, 2022), particularly to Split-Conformal (SC) prediction intervals, developed by Lei et al., (2018). Conformal prediction generates confidence regions with minimal assumptions, apart from the exchangeability of the data and symmetry of the regression function, and has correct finite-sample coverage by construction (see, e.g., Theorem 1 in Lei et al.,, 2018). However, conformal prediction requires training multiple estimators to create each confidence interval, leading to elevated computational costs in general. SC prediction is based on ranking non-conformity scores of exchangeable data, and reduces the computational burden of conformal prediction by splitting the data into two subsamples, one for fitting the model and another to rank the conformity scores, at the cost of halving the sample size in each task. Recently, Lugosi and Matabuena, (2024) modified the traditional SC ranking of residuals by estimating the quantiles of the errors, while Zhou and Müller, (2025) introduced Conditional Profile average transport Costs (CPCs) as a method to assess the compatibility of an element of the metric space with the conditional distribution of the response given a predictor value. The SC algorithm can be directly applied to develop confidence sets by using the conditional distribution of CPCs as a conformity score. However, the aforementioned references inherit the inefficiencies of SC inference derived from splitting the data into two subsamples.
@@
In this paper, we propose a novel method for constructing confidence regions for random objects in metric spaces using RF Out-Of-Bag (OOB) errors, which are a reliable estimate of predictive performance and are obtained automatically in the forest training process. Our approach builds on Zhang et al., (2020) by introducing OOB prediction balls, which are constructed around the RF prediction of the response with a radius determined by the $(1-\alpha)$-quantile of the empirical distribution of the OOB errors. For any given value of the response, a prediction ball can be generated with the information obtained when training a single forest, with OOB errors preventing any leakage of information about the response from the training phase. Although we derive OOB balls from RF predictions, they can be applied directly to any bagging regression model in metric spaces. We establish the asymptotic validity of OOB balls under four different coverage types: unconditional, and conditional on combinations of the training sample and prediction point. Our approach is related to SC inference, but is specifically tailored to RFs with metric-space-valued responses. By using OOB observations, OOB balls allow using the full data set for both prediction and uncertainty estimation, thereby achieving narrower confidence regions with improved coverage rates. We compare OOB and SC balls by simulating from a multivariate Euclidean linear model with different values for the dimensionality of the response, and conclude that the volume of OOB balls is smaller in general than that of SC balls, without compromising coverage. The coverage and radius of OOB balls are assessed using simulated data on the sphere, the hyperboloid, and the space of $2\times 2$ symmetric positive definite matrices. Additionally, we study the dynamics of sunspots on the Sun’s surface using a real data set. Specifically, we consider the problem of predicting the location of sunspots the last time they were observed, based on the first-ever observed location. The motion of sunspots is chaotic, so we use OOB balls to establish regions that contain the true location with a given probability. We use a data-driven distance on the sphere induced by the geodesic distance on the spheroid, illustrating the use of anisotropic distances to adapt the shape of OOB balls to the underlying distribution of the data.
@@ 2.1 Fréchet regression
Let $(\mathcal{Y},d_{\mathcal{Y}})$ denote the metric space where the response $Y$ is defined, with $d_{\mathcal{Y}}$ a distance function in $\mathcal{Y}$}. Denote by $(\mathcal{X},d_{\mathcal{X}}):=(\mathcal{X}_{1},d_{\mathcal{X}_{1}})\times\dots\times(\mathcal{X}_{p},d_{\mathcal{X}_{p}})$ the metric space of predictors, formed by the product of $p$ metric spaces. Analyzing metric space data requires basic statistical notions like the expectation and variance, which are undefined in the absence of a vector space structure. In order to generalize them for random objects, Fréchet, (1948) defined the homonymous mean and variance by exploiting the squared distance minimization property of the expectation in the Euclidean space:
@@
The absence of algebraic structure can hinder the generalization of Euclidean parametric regression models to metric spaces. Petersen and Müller, (2019) proposed a parametric model, known as global Fréchet regression, as an extension of multiple linear regression. They considered a totally bounded metric space $\mathcal{Y}$ and Euclidean predictors $\boldsymbol{X}\in\mathbb{R}^{p}$. Global Fréchet regression extends the linear regression model for Euclidean response, $m_{\mathrm{L}}(\boldsymbol{x}):=\beta_{0}+\boldsymbol{\beta}_{1}^{\top}(\boldsymbol{x}-\boldsymbol{\mu})$. Here, let $\boldsymbol{\mu}=\mathsf{E}(\boldsymbol{X})$ and $\boldsymbol{\Sigma}=\mathsf{Var}(\boldsymbol{X})$, and define the linear weight function $\omega(\boldsymbol{x}_{1},\boldsymbol{x}_{2}):=1+(\boldsymbol{x}_{1}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}_{2}-\boldsymbol{\mu}),$ $\boldsymbol{x}_{1},\,\boldsymbol{x}_{2}\in\mathbb{R}^{p}$. Petersen and Müller, (2019) showed that, when the response is Euclidean, the regression function is
@@
The empirical weights are defined as $n\omega_{i}(\boldsymbol{x}):=1+(\boldsymbol{X}_{i}-\bar{\boldsymbol{X}})^{\top}\widehat{\boldsymbol{\Sigma}}^{-1}(\boldsymbol{x}-\bar{\boldsymbol{X}}),$ with $\bar{\boldsymbol{X}}:=n^{-1}\sum_{i=1}^{n}\boldsymbol{X}_{i}$ and $\widehat{\boldsymbol{\Sigma}}:=n^{-1}\sum_{i=1}^{n}(\boldsymbol{X}_{i}-\bar{\boldsymbol{X}})(\boldsymbol{X}_{i}-\bar{\boldsymbol{X}})^{\top}.$ Petersen and Müller, (2019) derived pointwise and uniform consistency and convergence rates of the global estimator under several assumptions, which include the existence and uniqueness of $m_{\mathrm{L}}(\boldsymbol{x})$ and its estimator.
Global Fréchet regression is a parametric model that relies on the assumption of linearity. An alternative nonparametric approach to estimate the conditional Fréchet function $M_{\oplus}(x,y)$ in (2) is through a locally-weighted scheme of the form (3). Within this framework, Hein, (2009) proposed the Nadaraya–Watson estimator, considering weights $\omega_{i}$ defined through a kernel function $K$ and a bandwidth parameter $h$; that is, $\omega_{i}(x)=h^{-p}K(d_{\mathcal{X}}(X_{i},x)/h)$, where $p$ is the dimension of the metric space of predictors $(\mathcal{X},d_{\mathcal{X}})$. Under certain regularity conditions and when the predictors and the response belong to compact Riemannian manifolds, the kernel estimator is consistent.
Petersen and Müller, (2019) extended their global estimator to local Fréchet regression, a nonparametric approach of the form (3) that generalizes local linear regression (Fan and Gijbels,, 1996) for random objects and typically outperforms the Nadaraya–Watson estimator, especially near the boundaries of the support of the predictors. Petersen and Müller, (2019) derived the consistency of local Fréchet regression, and uniform rates of convergence were subsequently provided by Chen and Müller, (2022).
@@ 2.2 Random forests in metric spaces
Random forests (Breiman,, 2001) are among the most popular statistical learning techniques for prediction, and are particularly well-suited for extensions to metric spaces. Each decision tree begins with a root node containing all the training observations, which is recursively split into different nodes according to a splitting criterion until a stopping criterion is met. The nodes that are not split further are called terminal nodes. The prediction of the tree for a new observation $x$ is calculated as the average of training responses in the terminal node to which $x$ is assigned. The RF prediction is then computed as the average of the predictions from each tree of $x$. Two main adaptations are needed for RFs to apply to random objects: the splitting criterion and the aggregation of results to obtain predictions.
For a node $A\subset\mathcal{L}_{n}$ and a fixed split variable $X^{(j)}$, $j=1,\ldots,p$, Capitaine et al., (2024) proposed determining the left and right child nodes by the Voronoi partition of $A$ associated with the points $\left(c_{j,\ell},c_{j,r}\right)\in\mathcal{X}_{j}\times\mathcal{X}_{j}$. This partition divides $A$ based on distances from $c_{j,\ell}$ and $c_{j,r}$:
@@
Once the split is performed, the standard way to measure its quality is the CART (Classification And Regression Trees) criterion (Breiman et al.,, 1984). Capitaine et al., (2024) suggest an adaptation of this criterion for metric spaces, in terms of empirical Fréchet variances:
@@
For each variable $X^{(j)}$, in order to determine the splits $\left(c_{j,\ell},c_{j,r}\right)$, a split function is defined beforehand, which assigns a pair $(c_{j,\ell},c_{j,r})\in\mathcal{X}_{j}\times\mathcal{X}_{j}$ to any sample of $\mathcal{X}_{j}$. Capitaine et al., (2024) proposed the $2$-means algorithm as a flexible and computationally efficient choice for the split function. Once a split function is defined, the best possible split at node $A$ is the pair $\left(c_{j^{*},l},c_{j^{*},r}\right)$ that maximizes the decrease in Fréchet variance of the child nodes with respect to its parent node $A$, i.e., $j^{*}:=\operatorname*{arg\,max}_{j=1,\ldots,p}H_{j}\left(A,c_{j,\ell},c_{j,r}\right)$. In this way, observations are grouped into more homogeneous subgroups.
@@
The splits depend on the structure of the metric space of the predictors. For instance, if $\mathcal{X}=\mathbb{R}^{p}$, one could consider $(\mathcal{X},d_{\mathcal{X}})=(\mathbb{R},d_{2})\times\cdots\times(\mathbb{R},d_{2})$, but also $(\mathcal{X},d_{\mathcal{X}})=(\mathbb{R}^{p},d_{2})$. The first case connects with standard RFs, since the Fréchet mean and variance coincide with the mean and variance. In the second case, the data is split by a hyperplane $\{\boldsymbol{x}\in\mathbb{R}^{p}:d_{2}(\boldsymbol{x},\boldsymbol{c}_{l})=d_{2}(\boldsymbol{x},\boldsymbol{c}_{r})\}$, for a given pair $(\boldsymbol{c}_{l},\boldsymbol{c}_{r})\in\mathbb{R}^{p}\times\mathbb{R}^{p}$. This is conceptually related to oblique random forests (Li et al.,, 2023), where the split is given by the hyperplane that maximizes the decrease in variance.
@@
A similar idea allows the construction of RFs as an ensemble of decision trees in the metric case. This procedure, proposed by Capitaine et al., (2024), receives the name of Fréchet Random Forests (FRFs), which consist of a fixed number $B\in\mathbb{N}$ of decision trees, each built with a random resample $\mathcal{L}_{n}^{*b}$ of $\mathcal{L}_{n}$, $b=1,\ldots,B$. Commonly, nonparametric bootstrap resamples of size $n$ with replacement are employed. In addition, the best split in each non-terminal node of each tree is determined using a subset of $\ell\leq p$ features selected uniformly at random. The bagging process of subsampling from the training set stabilizes the variance of the trees that constitute the forest. In addition, random feature selection decreases the correlation between trees, generating different trees even if they were trained on the same subsample. Let $\widehat{m}_{T_{b}}(x,\boldsymbol{\theta}_{b})$ denote the prediction of the $b$-th tree for the point $x\in\mathcal{X}$ as in (5), where $\boldsymbol{\theta}_{b}$ is the randomization parameter. The prediction generated by the FRF at the point $x$ is defined as the Fréchet mean of the predictions of each of the trees:
@@
FRFs are applicable when both the predictors and the response are random objects. However, Capitaine et al., (2024) do not provide a theoretical analysis of the consistency of FRFs.
Qiu et al., (2024) refined and advanced the research conducted by Petersen and Müller, (2019) by developing a locally adaptive kernel for nonparametric Fréchet regression generated by Fréchet trees. Their approach, called Random Forest Weighted Local Constant Fréchet Regression (RFWLCFR), is equivalent to FRFs when $\mathcal{Y}=\mathbb{R}^{q}$, $q\in\mathbb{N}$, which can be seen from the property that the weighted sample mean minimizes the weighted sum of squared errors. Following the ideas of Meinshausen, (2006), the tree-building procedure in equation (5) naturally generates RF weights that can be used by an estimator of the form (3). Given a value $\boldsymbol{x}$ of the predictors in $\mathcal{X}=[0,1]^{p}$, a sample $\mathcal{L}_{n}$, and a bounded metric space $(\mathcal{Y},d_{\mathcal{Y}})$, the $b$-th tree provides the weights

### v2

@@ Abstract
Statistical methods for metric spaces provide a general and versatile framework for analyzing complex data types. We introduce a novel approach for constructing confidence regions around new predictions from any bagged regression algorithm with metric-space-valued responses. This includes the recent extensions of random forests for metric responses: Fréchet random forests (Capitaine et al., 2024), random forest weighted local constant Fréchet regression (Qiu et al., 2024), and metric random forests (Bulté and Sørensen, 2024). Our prediction regions leverage out-of-bag observations generated during a single forest training, employing the entire data set for both prediction and uncertainty quantification. We establish asymptotic guarantees of out-of-bag prediction balls for four coverage types under certain regularity conditions. Moreover, we demonstrate the superior stability and smaller radius of out-of-bag balls compared to split-conformal methods through extensive numerical experiments where the response lies on the Euclidean space, sphere, hyperboloid, and space of positive definite matrices. A real data application illustrates the potential of the confidence regions for quantifying the uncertainty in the study of solar dynamics and the use of data-driven non-isotropic distances on the sphere.
@@
In Section A, Theorems 4.1 and 4.2 and Corollaries 4.1 and 4.2 are proved, relying on four technical lemmas. Section B shows numerical experiments to validate the Fréchet means of Wishart-distributed random variables under the AI and LC distances, given in (15) and (16), respectively. Section C contains a comprehensive list of figures and tables from Section 5. In Section D, we reapproach the derivations by Jensen (1981, pp. 196–197) to clarify the connection of the HvMF and vMF distributions, and to detail the derivation and computation of the normalizing constant.
@@ 1 Introduction
As the complexity of data structures increases, traditional statistical methods that rely on vector space properties face significant limitations. Specialized methods have been developed for specific data types, including functional data (Li et al., 2022), directional data (Mardia and Jupp, 1999), compositional data (Aitchison, 1986), shape data (Dryden and Mardia, 2016), and covariance matrices (Dryden et al., 2009). Random variables from these complex data types are studied by Marron and Dryden (2021) under the terminology of random objects. A framework that has gained momentum in recent years is to consider random objects as elements in general metric spaces. While exploiting the structure of specific spaces can enhance the flexibility and tractability of statistical modeling, the generality of metric space analysis offers an unparalleled versatile framework that bypasses the need for algebraic structure or local Euclideanity, relying solely on the existence of a distance function.
@@
Key statistical concepts must be adapted to metric spaces. Fréchet (1948) generalized the mean through square distance minimization. Extensions of regression models for non-Euclidean data have primarily focused on responses defined in Riemannian manifolds (e.g., Fletcher, 2013; Kim et al., 2014). A foundational nonparametric regression estimator for metric output is the Nadaraya–Watson estimator, which uses local-constant weights determined by a kernel function and a bandwidth parameter. The Nadaraya–Watson estimator was extended by Hein (2009) to the case when both the response and the predictors are random objects using a (generalized) Fréchet mean. Building on the notion of Fréchet mean, Petersen and Müller (2019) introduced Fréchet regression, generalizing the standard Euclidean regression to random objects by estimating the conditional Fréchet mean. In the same work, global Fréchet regression was proposed as a parametric model that generalizes linear least squares regression for predictors in $\mathbb{R}^{p}$ and a metric response, alongside local Fréchet regression as a nonparametric generalization of local linear estimation for random object responses with Euclidean predictors.
@@
Random Forests (RFs; Breiman, 2001) are a popular nonparametric regression method widely recognized for their flexibility and predictive performance, owing to their ability to capture complex, nonlinear relationships, particularly in high-dimensional settings (Bühlmann and Yu, 2002; Hastie et al., 2009; Varian, 2014). In recent years, several authors have worked on adaptations of RFs to metric spaces under the framework of Fréchet regression. Capitaine et al. (2024) (first appeared as Capitaine et al., 2019) developed Fréchet Random Forests (FRFs), which adapt RFs for metric data by incorporating the Fréchet mean in the splitting criterion and the aggregation of trees. Alternatively, following the ideas of Meinshausen (2006), Qiu et al. (2024) leveraged the weights generated in the building of each tree to estimate the response as a weighted average of the training responses. Bulté and Sørensen (2024) introduced Fréchet medoids into the splitting criterion of the trees. Fréchet medoids serve as consistent estimators of the Fréchet mean and offer a computationally simpler alternative in settings where no closed-form expression is available and optimization over a manifold is typically employed.
Despite their predictive power, RFs have historically lacked well-developed tools for statistical inference, particularly for quantifying the uncertainty in a RF prediction. For Euclidean data, Meinshausen (2006) introduced quantile regression forests, which estimate conditional quantiles and naturally lead to prediction intervals. Barber et al. (2021) introduced jackknife+ to construct prediction intervals for regression, based on leave-one-out predictions and residuals. Kim et al. (2020) proposed jackknife+-after-bootstrap, which integrates ensemble learning with the jackknife+. Wang et al. (2023) give sufficient conditions for the robustness of RF predictions against small perturbations in the sample. They obtain stability bounds for the RF prediction, and use them to derive the coverage of RF prediction intervals.
@@
The construction of confidence regions in metric spaces has focused primarily on Split-Conformal (SC) prediction, developed by Lei et al. (2018). SC prediction is based on ranking non-conformity scores of exchangeable data, and reduces the computational burden of conformal prediction by splitting the data into two subsamples, one for fitting the model and another to rank the conformity scores, at the cost of halving the sample size in each task. Recently, Lugosi and Matabuena (2024) modified the traditional SC ranking of residuals by estimating the quantiles of the errors, while Zhou and Müller (2025) introduced Conditional Profile average transport Costs (CPCs) as a method to assess the compatibility of an element of the metric space with the conditional distribution of the response given a predictor value. The SC algorithm can be directly applied to develop confidence sets by using the conditional distribution of CPCs as a conformity score. However, the aforementioned references inherit the inefficiencies of SC inference derived from splitting the data into two subsamples.
@@
In this paper, we propose OOB prediction balls, a novel method for constructing confidence regions for random objects in metric spaces using RF Out-Of-Bag (OOB) errors, which are a reliable estimate of predictive performance and are obtained automatically in the forest training process. The OOB balls are constructed around the RF prediction of the response with a radius determined by the $(1-\alpha)$-quantile of the empirical distribution of the OOB errors. For any given value of the response, a prediction ball can be generated with the information obtained when training a single forest, with OOB errors preventing any leakage of information about the response from the training phase. Although we derive OOB balls from RF predictions, they can be applied directly to any bagging regression model in metric spaces. We establish the asymptotic validity of OOB balls under four different coverage types: unconditional, and conditional on combinations of the training sample and prediction point. Our approach is related to SC inference, but is specifically tailored to RFs with metric-space-valued responses. By using OOB observations, OOB balls allow using the full data set for both prediction and uncertainty estimation. We compare OOB and SC balls by simulating from a multivariate Euclidean linear model with different values for the dimensionality of the response, and conclude that the volume of OOB balls is smaller in general than that of SC balls, without compromising coverage. The coverage and radius of OOB balls are assessed using simulated data on the sphere, the hyperboloid, and the space of $2\times 2$ symmetric positive definite matrices. Additionally, we study the dynamics of sunspots on the Sun’s surface using a real data set. Specifically, we consider the problem of predicting the location of sunspots the last time they were observed, based on the first-ever observed location. The motion of sunspots is complex and difficult to predict, so we use OOB balls to establish regions that contain the true location with a given probability. We use a data-driven distance on the sphere induced by the geodesic distance on the spheroid, illustrating the use of anisotropic distances to adapt the shape of OOB balls to the underlying distribution of the data.
Our approach builds on Zhang et al. (2020), who proposed RF prediction intervals using out-of-bag observations and proved their asymptotic validity. The confidence regions of Zhang et al. (2020) have the shape of balls centered at the RF prediction and with radius given by a quantile of the empirical distribution of the OOB errors, which makes their generalization to metric spaces natural. Our method also relates to Wang et al. (2023), since it is based on the OOB residuals correctly approximating the RF prediction errors; and to Kim et al. (2020), when the ensemble learner is a bagged regressor. However, Wang et al. (2023) impose conditions on the tails of the square of the response variable, a condition that does not have a natural counterpart in general metric spaces. The approach of Kim et al. (2020) does not extend directly to general metric spaces, since the prediction intervals are built from empirical quantiles of the lower and upper interval endpoints, a step that relies on the order structure of the real line. While Kim et al. (2020) and Wang et al. (2023) provide non-asymptotic lower bounds for the coverage probability, these guarantees concern marginal coverage; our approach provides asymptotic guarantees for four different coverage types, as in Zhang et al. (2020). However, the absence of vector space structures complicates the proof of the coverage results.
@@ 2.1 Fréchet regression
Let $(\mathcal{Y},d_{\mathcal{Y}})$ denote the metric space where the response $Y$ is defined, with $d_{\mathcal{Y}}$ a distance function in $\mathcal{Y}$}. Denote by $(\mathcal{X},d_{\mathcal{X}}):=(\mathcal{X}_{1},d_{\mathcal{X}_{1}})\times\dots\times(\mathcal{X}_{p},d_{\mathcal{X}_{p}})$ the metric space of predictors, formed by the product of $p$ metric spaces. Analyzing metric space data requires basic statistical notions like the expectation and variance, which are undefined in the absence of a vector space structure. In order to generalize them for random objects, Fréchet (1948) defined the homonymous mean and variance by exploiting the squared distance minimization property of the expectation in the Euclidean space:
@@
The absence of algebraic structure can hinder the generalization of Euclidean parametric regression models to metric spaces. Petersen and Müller (2019) proposed a parametric model, known as global Fréchet regression, as an extension of multiple linear regression. They considered a totally bounded metric space $\mathcal{Y}$ and Euclidean predictors $\boldsymbol{X}\in\mathbb{R}^{p}$. Global Fréchet regression extends the linear regression model for Euclidean response, $m_{\mathrm{L}}(\boldsymbol{x}):=\beta_{0}+\boldsymbol{\beta}_{1}^{\top}(\boldsymbol{x}-\boldsymbol{\mu})$. Here, let $\boldsymbol{\mu}=\mathsf{E}(\boldsymbol{X})$ and $\boldsymbol{\Sigma}=\mathsf{Var}(\boldsymbol{X})$, and define the linear weight function $\omega(\boldsymbol{x}_{1},\boldsymbol{x}_{2}):=1+(\boldsymbol{x}_{1}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}_{2}-\boldsymbol{\mu}),$ $\boldsymbol{x}_{1},\,\boldsymbol{x}_{2}\in\mathbb{R}^{p}$. Petersen and Müller (2019) showed that, when the response is Euclidean, the regression function is
@@
The empirical weights are defined as $n\omega_{i}(\boldsymbol{x}):=1+(\boldsymbol{X}_{i}-\bar{\boldsymbol{X}})^{\top}\widehat{\boldsymbol{\Sigma}}^{-1}(\boldsymbol{x}-\bar{\boldsymbol{X}}),$ with $\bar{\boldsymbol{X}}:=n^{-1}\sum_{i=1}^{n}\boldsymbol{X}_{i}$ and $\widehat{\boldsymbol{\Sigma}}:=n^{-1}\sum_{i=1}^{n}(\boldsymbol{X}_{i}-\bar{\boldsymbol{X}})(\boldsymbol{X}_{i}-\bar{\boldsymbol{X}})^{\top}.$ Petersen and Müller (2019) derived pointwise and uniform consistency and convergence rates of the global estimator under several assumptions, which include the existence and uniqueness of $m_{\mathrm{L}}(\boldsymbol{x})$ and its estimator.
Global Fréchet regression is a parametric model that relies on the assumption of linearity. An alternative nonparametric approach to estimate the conditional Fréchet function $M_{\oplus}(x,y)$ in (2) is through a locally-weighted scheme of the form (3). Within this framework, Hein (2009) proposed the Nadaraya–Watson estimator, considering weights $\omega_{i}$ defined through a kernel function $K$ and a bandwidth parameter $h$; that is, $\omega_{i}(x)=h^{-p}K(d_{\mathcal{X}}(X_{i},x)/h)$, where $p$ is the dimension of the metric space of predictors $(\mathcal{X},d_{\mathcal{X}})$. Under certain regularity conditions and when the predictors and the response belong to compact Riemannian manifolds, the kernel estimator is consistent.
Petersen and Müller (2019) extended their global estimator to local Fréchet regression, a nonparametric approach of the form (3) that generalizes local linear regression (Fan and Gijbels, 1996) for random objects and typically outperforms the Nadaraya–Watson estimator, especially near the boundaries of the support of the predictors. Petersen and Müller (2019) derived the consistency of local Fréchet regression, and uniform rates of convergence were subsequently provided by Chen and Müller (2022).
@@ 2.2 Random forests in metric spaces
Random forests (Breiman, 2001) are among the most popular statistical learning techniques for prediction, and are particularly well-suited for extensions to metric spaces. Each decision tree begins with a root node containing all the training observations, which is recursively split into different nodes according to a splitting criterion until a stopping criterion is met. The nodes that are not split further are called terminal nodes. The prediction of the tree for a new observation $x$ is calculated as the average of training responses in the terminal node to which $x$ is assigned. The RF prediction is then computed as the average of the predictions from each tree of $x$. Two main adaptations are needed for RFs to apply to random objects: the splitting criterion and the aggregation of results to obtain predictions.
For a node $A\subset\mathcal{L}_{n}$ and a fixed split variable $X^{(j)}$, $j=1,\ldots,p$, Capitaine et al. (2024) proposed determining the left and right child nodes by the Voronoi partition of $A$ associated with the points $\left(c_{j,\ell},c_{j,r}\right)\in\mathcal{X}_{j}\times\mathcal{X}_{j}$. This partition divides $A$ based on distances from $c_{j,\ell}$ and $c_{j,r}$:
@@
Once the split is performed, the standard way to measure its quality is the CART (Classification And Regression Trees) criterion (Breiman et al., 1984). Capitaine et al. (2024) suggest an adaptation of this criterion for metric spaces, in terms of empirical Fréchet variances:
@@
For each variable $X^{(j)}$, in order to determine the splits $\left(c_{j,\ell},c_{j,r}\right)$, a split function is defined beforehand, which assigns a pair $(c_{j,\ell},c_{j,r})\in\mathcal{X}_{j}\times\mathcal{X}_{j}$ to any sample of $\mathcal{X}_{j}$. Capitaine et al. (2024) proposed the $2$-means algorithm as a flexible and computationally efficient choice for the split function. Once a split function is defined, the best possible split at node $A$ is the pair $\left(c_{j^{*},l},c_{j^{*},r}\right)$ that maximizes the decrease in Fréchet variance of the child nodes with respect to its parent node $A$, i.e., $j^{*}:=\operatorname*{arg\,max}_{j=1,\ldots,p}H_{j}\left(A,c_{j,\ell},c_{j,r}\right)$. In this way, observations are grouped into more homogeneous subgroups.
@@
The splits depend on the structure of the metric space of the predictors. For instance, if $\mathcal{X}=\mathbb{R}^{p}$, one could consider $(\mathcal{X},d_{\mathcal{X}})=(\mathbb{R},d_{2})\times\cdots\times(\mathbb{R},d_{2})$, but also $(\mathcal{X},d_{\mathcal{X}})=(\mathbb{R}^{p},d_{2})$. The first case connects with standard RFs, since the Fréchet mean and variance coincide with the mean and variance. In the second case, the data is split by a hyperplane $\{\boldsymbol{x}\in\mathbb{R}^{p}:d_{2}(\boldsymbol{x},\boldsymbol{c}_{l})=d_{2}(\boldsymbol{x},\boldsymbol{c}_{r})\}$, for a given pair $(\boldsymbol{c}_{l},\boldsymbol{c}_{r})\in\mathbb{R}^{p}\times\mathbb{R}^{p}$. This is conceptually related to oblique random forests (Li et al., 2023), where the split is given by the hyperplane that maximizes the decrease in variance.
@@
A similar idea allows the construction of RFs as an ensemble of decision trees in the metric case. This procedure, proposed by Capitaine et al. (2024), receives the name of Fréchet Random Forests (FRFs), which consist of a fixed number $B\in\mathbb{N}$ of decision trees, each built with a random resample $\mathcal{L}_{n}^{*b}$ of $\mathcal{L}_{n}$, $b=1,\ldots,B$. Commonly, nonparametric bootstrap resamples of size $n$ with replacement are employed. In addition, the best split in each non-terminal node of each tree is determined using a subset of $\ell\leq p$ features selected uniformly at random. The bagging process of subsampling from the training set stabilizes the variance of the trees that constitute the forest. In addition, random feature selection decreases the correlation between trees, generating different trees even if they were trained on the same subsample. Let $\widehat{m}_{T_{b}}(x,\boldsymbol{\theta}_{b})$ denote the prediction of the $b$-th tree for the point $x\in\mathcal{X}$ as in (5), where $\boldsymbol{\theta}_{b}$ is the randomization parameter. The prediction generated by the FRF at the point $x$ is defined as the Fréchet mean of the predictions of each of the trees:
@@
FRFs are applicable when both the predictors and the response are random objects. However, Capitaine et al. (2024) do not provide a theoretical analysis of the consistency of FRFs.
Qiu et al. (2024) refined and advanced the research conducted by Petersen and Müller (2019) by developing a locally adaptive kernel for nonparametric Fréchet regression generated by Fréchet trees. Their approach, called Random Forest Weighted Local Constant Fréchet Regression (RFWLCFR), is equivalent to FRFs when $\mathcal{Y}=\mathbb{R}^{q}$, $q\in\mathbb{N}$, which can be seen from the property that the weighted sample mean minimizes the weighted sum of squared errors. Following the ideas of Meinshausen (2006), the tree-building procedure in equation (5) naturally generates RF weights that can be used by an estimator of the form (3). Given a value $\boldsymbol{x}$ of the predictors in $\mathcal{X}=[0,1]^{p}$, a sample $\mathcal{L}_{n}$, and a bounded metric space $(\mathcal{Y},d_{\mathcal{Y}})$, the $b$-th tree provides the weights